a large boulder is ejected vertically upward from a volcano with an initial speed of 40.0m/sec. Air resistance may be ignored

Grade:11

1 Answers

Arun
6 years ago
Dear Student
your question is in complete.
but I have solved the question somewhere so here I am again solving
A large boulder is ejected vertically upward from a volcano with an initial speed of 40.0 m/s. Air resistance may be ignored. 

(a) At what time after being ejected is the boulder moving at 20.0 m/s upward? 

(b) At what time is it moving at 20.0 m/s down-ward? 

(c) When is the displacement of the boulder from its initial position zero? 
Solution
 
 
(a) V = Vo + at 
20 = 40 - 9.8*t → t = 2.04 s 

(b) -20 = 40 - 9.8*t → t = 6.12 s 

(c) s = 0 = Vo*t + ½at² = (Vo + ½at) * t 
So one solution is obviously t = 0 s. The other is 
0 = 40 - ½ * 9.8m/s² * t → t = 8.16 s 
== OR == 
find out when its velocity is -40 m/s: 
-40 = 40 - 9.8*t
 
Regards
Arun (askIITians forum expert)